OpenStudy (anonymous):

sin(20-π)= sinΘ SOLVE ON THE INTERVAL 0≤Θ<2π

OpenStudy (mary.rojas):

@shamil98 @inkyvoyd @satellite73 @timo86m ???sos

OpenStudy (anonymous):

anyone help?

OpenStudy (mary.rojas):

@Mertsj SOS??

OpenStudy (mary.rojas):

Dont give up lol

OpenStudy (mertsj):

Is that 20-pi?

OpenStudy (mertsj):

\[\sin(20-\pi)=-\sin 20\]

OpenStudy (mertsj):

\[-\sin 20=\sin \theta \]

OpenStudy (mertsj):

And I don't know what you mean by saying that the domain is 0 to 0

OpenStudy (anonymous):

0 to 2π i meant

OpenStudy (anonymous):

help? is that right?